OpenStudy (lukecrayonz):

Please lord help, I have a test tomorrow & I do not understand this one thing. 1. 7^(2x+1)=3^(x+3) 2. 9^(x+2)=2^(5x-4) 3. 3^(4x+3)=8^(-x+2) Please lord help, I have a test tomorrow & I do not understand this one thing. 1. 7^(2x+1)=3^(x+3) 2. 9^(x+2)=2^(5x-4) 3. 3^(4x+3)=8^(-x+2) @Mathematics

OpenStudy (lukecrayonz):

You're solving for x :O So yeah. I'll repost this in the morning if nobody knows right now.

hero (hero):

take logs of both sides

OpenStudy (lukecrayonz):

Hero, my online calculator (not wolfram) shows to use ln? and they got the right answer according to my textbook.

hero (hero):

(2x+1)log(7) = (x+3)log(3)

hero (hero):

\[\frac{2x+1}{x+3} = \frac{\log(3)}{\log{7}}\]

hero (hero):

simplify right side

hero (hero):

then continue solving for x

OpenStudy (lukecrayonz):

My textbook, shows for its answer: ln(27/7) over ln(49/3), which equals x=.48

hero (hero):

You can use ln or logs...doesn't matter

hero (hero):

You will get the same answer

hero (hero):

But use ln since your book says to use it

hero (hero):

The procedure is still the same

OpenStudy (lukecrayonz):

I'm still lost. Can you do number two?

hero (hero):

When you take logs, whatever is in the exponent comes down from being an exponent and then all you have is just the log of the base of the exponent

OpenStudy (lukecrayonz):

I get all that, I just don't get how you actually solved to get x=.48 from that problem.

hero (hero):

9^(x+2)=2^(5x-4) Take logs of both sides: (x+2)log(9) = (5x-4) log(2) Divide both sides by (5x-4), then by log(9): (x+2)/(5x-4) = log(2)/log(9)

hero (hero):

I thought you wanted help with the next problem

hero (hero):

Okay, I'll finish solving the first problem

OpenStudy (lukecrayonz):

I do but i thought you were stil concerntrating on the first. just keep going with #2

hero (hero):

First of all, if you're trying to do this without a calculator, not a good idea.

hero (hero):

You need a calculator to evaluate log(3) and log(7)

OpenStudy (lukecrayonz):

0.56

hero (hero):

Yes, now multiply both sides by (x+3) again

hero (hero):

2x+1 = .56(x+3)

hero (hero):

distribute the .56

OpenStudy (lukecrayonz):

.56x+1.68

hero (hero):

2x+1 = .56x + 1.6973

OpenStudy (lukecrayonz):

And the answer is .48! :D

hero (hero):

:D

hero (hero):

I thought the answer was supposed to be 48

hero (hero):

Oh nevermind, lol

hero (hero):

You got it

hero (hero):

It wasn't that hard man

OpenStudy (lukecrayonz):

Im looking at the 2nd one, and I think i am doing something wrong..

hero (hero):

It's the same exact procedure man

OpenStudy (lukecrayonz):

(x+2)/(5x-4) = log(2)/log(9) .31

hero (hero):

Also, you can verify your answer by plugging x back in

OpenStudy (lukecrayonz):

K so i distribute .31 through 5x-4

hero (hero):

and make sure both sides are the same

hero (hero):

on the test, plug x back in and verify!

OpenStudy (lukecrayonz):

x+2=-1.24+1.55 x

hero (hero):

That's the only way to verify that you're right

hero (hero):

If you don't do that, you will fail

hero (hero):

If you forget to do it, you will fail

OpenStudy (lukecrayonz):

x+2=-1.24+1.55 x x=5.89091

hero (hero):

I can promise you that

hero (hero):

So make sure to verify your answer tomorrow

OpenStudy (lukecrayonz):

I know, thats what I was planning on doing the whole test, just checking constantly.

hero (hero):

Good

hero (hero):

I just wanted you to confirm that you will do it

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

hero (hero):

Do not round answers to the end. Never ever round an answer until the end

hero (hero):

That's rule #1 of how not to make a mistake on a math test

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

Eh, my teacher is lenient and doesn't care as long as you are somewhere in the ballpark.

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

Like with a calculator I don't round because I just do 2nd answer, and just input like 50 decimals :P

hero (hero):

You're lucky. 99% of most teachers aren't so nice. But still, you want to get as close to the right answer as possible. I wouldn't make it a habit of rounding prematurely man

hero (hero):

Don't ever take an online exam where they want the exact answer

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

Do you know ANY way of memorzing formulas besides writing them a million times?

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

I was thinking of just learning the formulas like before the test, and right when the test starts just write them right on the test..

hero (hero):

Yep, you just have to know them man. That's all I can say. Most of it is just algebra anyway

hero (hero):

What is the test on?

hero (hero):

What chapter or section of math topics does it cover?

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

Exponential functions, Logarithmic functions, Properties of Logarithms, Exponential and Logarithmic Equations, and inverses of logarithms/exponential equations

hero (hero):

log(a-b) + log(a-b) = ?

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

inverse as in y=x+2, then just replacing x for y, y for x, solving for y. x=y+2, y=x-2, easy.

hero (hero):

Well, actually, that's a bad one because the answer to that would just be 2(log(a-b)) which if you further simplify to 2(log(a/b))

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

I have no idea what that is :S

hero (hero):

you see that they have the same base so you combine them

hero (hero):

It's like x + x

hero (hero):

= 2x

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

oh! I know what that is, i thought you wanted me to name something lmao

hero (hero):

except it has an extra step included

OpenStudy (lukecrayonz):

Like as if its a property or something, i was like uhhh....

hero (hero):

Yeah , but you didn't get it immediately

hero (hero):

No, I want you to simplify it

hero (hero):

c'mon man, you should know what to do if someone gave you that

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

You multiply them btw ;)

hero (hero):

You can't put "oh yeah I know what that is" on the test

hero (hero):

If they're the same base you can just add and simplify as I showed above

hero (hero):

You're forgetting the x + x rule

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

ln(x+2)+ln(3x-2)=2ln(2x), you do ln(x+2)(3x-2)=ln4x^2, 3x^2+4x-4=4x^2 x^2-4x+4=0 quadraticccc

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

yeah my teacher sucks she doesnt teach us rules, she just tells us to solve.

hero (hero):

ln(3x-2)+ln(3x-2) = ?

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

(3x-2)(3x-2) 9x^2-6x-6x+4 9x^2-12x+4

OpenStudy (lukecrayonz):

Textbook got 5.65, but they round weird, so 5.89 is close enough :D

OpenStudy (lukecrayonz):

or 2ln(3x-2)